# Latent space models for grouped multiplex networks

### v1

@@ Appendix B Proofs
Here, we present proofs of all the statements made throughout this manuscript.
B.1 Section 2.2 proofs
@@
By the first condition and Lemma 1 in (multiness), for any $k=1,\ldots,K$ and $\ell=1,\ldots,m_{k}$, we have
@@
Consider the layers $(k_{1},\ell_{1})$ and $(k_{2},\ell_{2})$ satisfying the third identifiability condition. Writing (39) for them and equating two expressions for $V^{\prime}$, we obtain for every $r=1,\ldots,d_{0}$
@@
By linear independence, it implies that $o_{k_{1}\ell_{1},r}^{(2)}=o_{k_{1}\ell_{1},r}^{(3)}=0$. Since it holds for any $r=1,\ldots,d_{0}$, we have $O_{21,k_{1}\ell_{1}}=O_{31,k_{1}\ell_{1}}=0$. By (39), it means $V^{\prime}=VO_{k_{1},\ell_{1}}$ and so for any layer $(k,\ell)$, we have $O_{21,k\ell}=O_{31,k\ell}=0$. Using the fact that $S_{k\ell}O_{k\ell}S_{k\ell}^{\top}$ is an indefinite orthogonal transformation, we can derive that the symmetric upper-diagonal blocks $O_{12,k\ell},O_{13,k\ell}$ are also zeros.
Now, consider layers $s$ and $t$ in group $k$ satisfying the second identifiability condition. Writing (39) for them and equating two expressions for $W_{k}^{\prime}$, we have for every $r=d_{0}+1,\ldots,d_{0}+d_{k}$
@@
By linear independence, it implies that $o_{ks,r}^{(3)}=o_{kt,r}^{(3)}=0$. Since it holds for any $r=d_{0}+1,\ldots,d_{0}+d_{k}$, we deduce $O_{32,ks}=0$. Combining with the previously established $O_{12,ks}=0$ this implies that $W_{k}^{\prime}=W_{k}O_{22,ks}$ and so for any layer $(k,\ell)$, we have $O_{32,k\ell}=0$. Since the lower right $2\times 2$ block sub-matrix of $O_{k\ell}$ is also an indefinite orthogonal rotation (up-to conjugation with a permutation matrix), we deduce $O_{23,k\ell}=0$, which completes the proof.
B.2 Section 3 Proofs
@@
Under the Gaussian distribution, for each $1\leq i\leq j\leq n$, we can rewrite the joint negative log-likelihood of the $(i,j)$-th entries (up to $1/2\sigma^{2}$ multiplier) as
$\sum_{(k,\ell)\in\mathcal{I}}(A_{k\ell,ij}-S_{ij}-Q_{k,ij}-\hat{R}_{k\ell,ij})^{2}$ $=\sum_{k=1}^{K}\Bigl[m_{k}\bigl(S_{ij}+Q_{k,ij}\bigr)^{2}-2m_{k}\tilde{A}_{k,ij}(S_{ij}+Q_{k,ij})\Bigr]+const$
$=\sum_{k=1}^{K}m_{k}\bigl(S_{ij}+Q_{k,ij}-\tilde{A}_{k,ij}\bigr)^{2}+const,$
where constconst denotes terms independent of $S$ and $Q_{k}$. Therefore, the solution to Problem (11) coincides with the solution of the following program:
@@
With the assumed edge-entry distribution, Problem (10) solves
@@
where the off-diagonal entries of each matrix inside the Frobenius norm are counted twice compared to (9), leading to an additional $1/2$ multiplier in front of the sum. By optimality, $\widehat{S+Q_{k}}$ should minimize the target function in (40) over $S+Q_{k}$ with each $R_{k\ell}$ fixed to $\hat{R}_{k\ell}$, that is, solve the optimization problem
@@
where constconst denotes terms independent of $Z$. By the variational definition of soft-thresholding, the optimal $Z$ in the last problem is achieved at (21). ∎
B.3 Section 4 Proofs
Throughout this section, we will use the following additional notations. For $a,b\in\mathbb{R}$, denote $a\wedge b=\min(a,b),\ a\vee b=\max(a,b)$. For a matrix $V\in\mathbb{R}^{n\times d}$, denote the projector on its column space by $\mathcal{P}_{V}=V(V^{\top}V)^{\dagger}V^{\top}$ where $(\cdot)^{\dagger}$ is the Moore-Penrose pseudoinverse. For a symmetric matrix $Z\in\mathbb{R}^{n\times n}$ with the eigen-decomposition $Z=V\Gamma V^{\top}$, we will assume by default that the eigenvalues in $\Gamma=\operatorname{diag}[\gamma_{1}(Z),\ldots,\gamma_{n}(Z)]$ are sorted in descending order wrt their absolute value:
$\|\Gamma\|_{2}=|\gamma_{1}(Z)|\geq\ldots\geq|\gamma_{n}(Z)|$
where the smallest (in absolute value) nonzero eigenvalue of $Z$ is denoted by $\gamma_{\min}(Z)$.
In this section, we will use multiple measures of distance between subspaces spanned by the columns of two $n\times d$ orthogonal matrices $V$ and $\hat{V}$. Suppose the singular values of $V^{\top}\hat{V}$ are $\sigma_{1}\geq$ $\sigma_{2}\geq\cdots\geq\sigma_{d}\geq 0$. Then, following standard notations, we define the $\sin\Theta$-matrix as
$\sin\Theta(V,\hat{V})=\operatorname{diag}\left[\sin(\cos^{-1}\left(\sigma_{1}\right)),\cdots,\sin(\cos^{-1}\left(\sigma_{d}\right))\right].$
and use $\|\sin\Theta(V,\hat{V})\|_{2}$ to measure the distance. For symmetric matrices $Z,\hat{Z}\in\mathbb{R}^{n\times n}$ with top $d<n$ eigenvectors denoted respectively as $V,\hat{V}\in\mathbb{R}^{n\times d}$, it would be convenient to define:
$\sin_{d}(Z,\hat{Z}):=\|\sin\Theta(V,\hat{V})\|_{2}.$
Notice that $V$ and $\hat{V}$ (and thus the defined quantity above) are uniquely defined only if there is a gap between the $d$-th and $(d+1)$-st eigenvalues of $Z$ and $\hat{Z}$ or these two eigenvalues are zero. In all further applications, this will be satisfied for sufficiently large $n$, as $d$ will denote the rank of $Z$ and $\hat{Z}=\mathcal{T}_{\rho}(Z+E)$ will be a soft-thresholded perturbation of $Z$ with an additive noise matrix $E$ satisfying $\|E\|_{2}=o(\gamma_{\min}(Z))$.
Lemma 1 in cai_zhang establishes the equivalence between the sin distance and several other natural metrics. Here, we only state several useful corollaries from this Lemma that we will further need throughout this section:
@@ B.3 Section 4 Proofs
Another important result we will need further is Theorem 1 of cai_zhang, which provides a useful bound on the $\sin$-distance between the eigenspaces of a matrix and its perturbation. Since originally this result was formulated for arbitrary matrices and in this work, we only deal with symmetric matrices, we restate it in Lemma B.2 with less notational load and a slightly loosened but much more algebraically convenient bound.
We start by giving an outline of the proof for Theorem 4.1. The initial step is to rewrite the first iteration update for each parameter in Remark 5 as a soft-thresholding operator applied to the parameter’s ground-truth value, say $Z$, plus the error term, which comprises Gaussian noise and the weighted sum of an update-dependent set of latent components errors:
@@
For example, the first stage update rule for the parameters $S+Q_{k}$ and $\{R_{k\ell}\}_{\ell=1}^{m_{k}}$ can be rewritten
@@
where by $\Delta_{Z}^{(t)}:=Z^{(t)}-Z$ we denote the error of parameter $Z$ after $t$ iterations. The second stage update rule at the first iteration can be written similarly:
$Q_{k}^{(1)}=\mathcal{T}_{\rho_{1k}}\Bigl[Q_{k}-\Delta_{S}^{(0)}-{1\over m_{k}}\sum_{\ell=1}^{m_{k}}\Delta_{R_{k\ell}}+\bar{E}_{k}\Bigr],\qquad k=1,\ldots,K,$ (45)
$S^{(1)}=\mathcal{T}_{\rho_{0}}\Bigl[S-\sum_{k=1}^{K}{m_{k}\over M}\Delta_{Q_{k}}^{(1)}-{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Delta_{R_{k\ell}}+\bar{E}\Bigr],$
where $\Delta_{R_{k\ell}},(k,\ell)\in\mathcal{I}$ denotes the individual component error after the last iteration of the $k$-th group updates in the first stage. We do not consider further ($t>1$) iterations in the update procedures of both stages, since in the theoretical framework we use, it is impossible to demonstrate the rate improvement after more than one iteration update. However, we can demonstrate that the error would not increase after additional iterations. Thus, by induction, the final error rate obtained by running the updates until numerical convergence will match the rate achieved after the first iteration.
The key to our proof will be Lemma B.3 that shows that the soft threshold operator applied to a matrix perturbed by additive error as in (43) should have the threshold $\rho$ with the rate of the total error’s spectral norm to guarantee that the estimate $\hat{Z}$ is close to the ground truth $Z$ in the Frobenius norm. In particular, it should dominate both the spectral norm of the other components’ errors, which we control using Lemma B.4, and the average of the appropriate Gaussian noise matrices, which we control by restricting the analysis to the set $\mathcal{E}_{noise}$ in (26). Importantly, the latter event occurs with an overwhelming probability according to Lemma B.1.
We postpone the proof of the mentioned technical lemmas to Section B.4 and give a formal proof of Theorem 4.1.
Proof of Theorem 4.1.
Assume that the events $\mathcal{E}_{noise}$ in (26) and $\mathcal{E}_{init}$ in (28) occur simultaneously. By Lemma B.4 and Assumption 3, probability of that is at least $1-(M+K+1)ne^{-C_{0}n}$ as needed.
First stage (group $k$): We first establish the properties of $R_{k\ell}^{(1)},\ \ell=1,\ldots,m_{k}$ by applying Lemma B.3 with the following correspondence of notations due to (44)
@@
and by Property 4, if $R_{k\ell}$ is PSD, then $R_{k\ell}^{(1)}$ is also PSD. Combining Corollary B.1 and (46), we have for sufficiently large $n$ that
@@
which implies together with (46) that for sufficiently large $n$ it holds
@@
Now, let us establish an error bound for $\Delta_{S+Q_{k}}^{(1)}$. With $C_{2}:={11(1+3\sigma)B_{R}\over b_{R}(1\wedge b_{R})}$, define
$\rho_{1k}=C_{2}\max_{1\leq\ell\leq m_{k}}\rho_{1k\ell}\Bigl[{1\over m_{k}}\vee s_{u,u}^{(k)}\vee n^{-\tau}\max_{1\leq\ell\leq m_{k}}\rho_{1k\ell}\Bigr]^{1/2}$ (50)
so that it holds due to $\rho_{1k\ell}\geq 3\sigma(n/m_{k})^{1/2}$
@@
Then, Properties 1 and 2 in Lemma B.3 imply
@@
Combining (50) and (52), we can see that $\|\Delta_{S+Q_{k}}^{(1)}\|_{2}=o(r_{k}^{(I)}\vee n^{1/2})$. So, treating $(S+Q_{k})^{(1)}$ as a new initializer for $S+Q_{k}$, we can see that all bounds dependent on it, that is, (46) and (48), are still satisfied. Thus, by induction, the next iteration errors $\Delta_{R_{k\ell}}^{(t)}$ and $\Delta_{S+Q_{k}}^{(t)}$ for any $t\geq 1$ satisfy the error bounds in (47) and (52), respectively.
Second stage: We first establish the properties of $Q_{k}^{(1)},\ k=1,\ldots,K$ by applying Lemma B.3 with the following correspondence of notations due to (45)
@@
So, Lemma B.4 with $s:=s_{w,w}$, $\theta:=\max_{1\leq k\leq K}\theta_{k}$, and $\rho:=\max_{1\leq k\leq K}\rho_{2k}$, we have for sufficiently large $n$ such that $\pi_{\max}\geq Km_{k}/M$ (guaranteed by Assumption 1):
$\Bigl\|{1\over M}\sum_{k=1}^{K}m_{k}\Delta_{Q_{k}}^{(1)}\Bigr\|_{2}\leq\pi_{\max}\Bigl\|{1\over K}\sum_{k=1}^{K}\Delta_{Q_{k}}^{(1)}\Bigr\|_{2}\leq{11B_{Q}\pi_{\max}\over b_{Q}(1\wedge b_{Q})}\max_{1\leq k\leq K}\rho_{2k}\Bigl[{1\over K}\vee s_{w,w}\vee n^{-\tau}\max_{1\leq k\leq K}\rho_{2k}\Bigr]^{1/2}.$
Now, let us establish an error bound for $\Delta_{S}^{(1)}$. Notice that (49) holds for the average of any iteration errors $\Delta_{R_{k\ell}}^{(t)}$, and thus also for the average of errors $\Delta_{R_{k\ell}}$ output at the end of the first stage. Thus, by Lemma B.4 with $s:=s_{u,u},\theta:=\max_{(k,\ell)\in\mathcal{I}}\theta_{k\ell}$, and $\rho:=\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}$:
$\Bigl\|{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Delta_{R_{k\ell}}\Bigr\|_{2}$ $\leq 11{B_{R}\over b_{R}}\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}\Bigr[{1\over M}\vee s_{u,u}\vee{1\over b_{R}n^{\tau}}\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}\Bigl]^{1/2}$
With $C_{3}:=12\bigl({B_{Q}\pi_{\max}\over b_{Q}(1\wedge b_{Q})}+{B_{R}\over b_{R}(1\wedge b_{R})}\bigr)$, define
@@
so that due to ${\rho_{1k\ell}/M^{1/2}}\geq 3\sigma(n/M)^{1/2}\geq\|\bar{E}\|_{2}$ it holds:
$\Bigl\|{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Delta_{R_{k\ell}}^{(1)}\Bigr\|_{2}+\Bigl\|{1\over M}\sum_{k=1}^{K}m_{k}\Delta_{Q_{k}}^{(1)}\Bigr\|_{2}+\|\bar{E}\|_{2}<\rho_{2}$
@@
and by Property 4, if $S$ is PSD, then $S^{(1)}$ is also PSD. ∎

### v2

@@ Appendix B Proofs
B.1 Section 2.2 proofs (Identifiability)
@@
By condition 1 and Lemma 1 in (MacDonald et al., 2022), for any $k=1,\ldots,K$ and $\ell=1,\ldots,m_{k}$, we have
@@
Consider layers $(k_{1},\ell_{1})$ and $(k_{2},\ell_{2})$ satisfying condition 3. Plugging them into (30) and equating two expressions for $V^{\prime}$, we obtain, for every $r=1,\ldots,d_{0}$
@@
By linear independence, it implies that $o_{k_{1}\ell_{1},r}^{(2)}=o_{k_{1}\ell_{1},r}^{(3)}=0$. Since it holds for any $r=1,\ldots,d_{0}$, we have $O_{21,k_{1}\ell_{1}}=O_{31,k_{1}\ell_{1}}=0$. By (30), it means $V^{\prime}=VO_{k_{1},\ell_{1}}$ and so for any layer $(k,\ell)$, we have $O_{21,k\ell}=O_{31,k\ell}=0$. Using the fact that $S_{k\ell}O_{k\ell}S_{k\ell}^{\top}$ is an indefinite orthogonal transformation, we can conclude that the symmetric upper-diagonal blocks $O_{12,k\ell},O_{13,k\ell}$ are also zeros.
Now, consider layers $s$ and $t$ in group $k$ satisfying the second identifiability condition. Plugging them into (30) and equating two expressions for $W_{k}^{\prime}$ now, we have, for every $r=d_{0}+1,\ldots,d_{0}+d_{k}$
@@
By linear independence, $o_{ks,r}^{(3)}=o_{kt,r}^{(3)}=0$. Since this holds for all $r=d_{0}+1,\ldots,d_{0}+d_{k}$, we deduce $O_{32,ks}=0$. Combined with $O_{12,ks}=0$, this implies that $W_{k}^{\prime}=W_{k}O_{22,ks}$ and so for any layer $(k,\ell)$, we have $O_{32,k\ell}=0$. Since the lower-right $2\times 2$ block sub-matrix of $O_{k\ell}$ is also an indefinite orthogonal rotation, we deduce $O_{23,k\ell}=0$, which completes the proof. ∎
B.2 Section 3 Proofs (Optimization)
@@
For the Gaussian distribution, for each $1\leq i\leq j\leq n$, we can rewrite the joint negative log-likelihood of the $(i,j)$-th entries (up to the $1/2\sigma^{2}$ multiplier) as
$\sum_{(k,\ell)\in\mathcal{I}}(A_{k\ell,ij}-S_{ij}-Q_{k,ij}-\hat{R}_{k\ell,ij})^{2}$ $=\sum_{k=1}^{K}m_{k}\bigl(S_{ij}+Q_{k,ij}-\tilde{A}_{k,ij}\bigr)^{2}+const,$
where constconst is a term independent of $S$ and $Q_{k}$. Therefore, the solution to Problem (11) coincides with the solution of
@@
With the assumed edge distribution, Problem (10) solves
@@
where the off-diagonal entries of each matrix inside the Frobenius norm are counted twice compared to (9), leading to an additional $1/2$ multiplier in front of the sum. By optimality, $\widehat{S+Q_{k}}$ should minimize the target function in (31) over $S+Q_{k}$ with each $R_{k\ell}$ fixed to $\hat{R}_{k\ell}$, that is,
@@
where constconst is a term independent of $Z$. By the variational definition of soft-thresholding, the optimal $Z$ in the last problem is achieved at (18). ∎
B.3 Section 4 Proofs (Consistency)
We start with additional notation. For $a,b\in\mathbb{R}$, let $a\wedge b=\min(a,b),\ a\vee b=\max(a,b)$. For a matrix $V\in\mathbb{R}^{n\times d}$, denote the projection operator onto its column space by $\mathcal{P}_{V}=V(V^{\top}V)^{\dagger}V^{\top}$, where $(\cdot)^{\dagger}$ is the Moore-Penrose pseudoinverse. For a symmetric matrix $Z\in\mathbb{R}^{n\times n}$ with the eigendecomposition $Z=V\Gamma V^{\top}$, we always order the diagonal entries of $\Gamma$ in descending order by their absolute values, with
$\|\Gamma\|_{2}=|\gamma_{1}(Z)|\geq\ldots\geq|\gamma_{n}(Z)|=\gamma_{\min}(Z).$
In this section, we use multiple measures of similarity between subspaces spanned by the columns of two $n\times d$ orthogonal matrices $V$ and $\hat{V}$. Suppose the singular values of $V^{\top}\hat{V}$ are $\sigma_{1}\geq$ $\sigma_{2}\geq\cdots\geq\sigma_{d}\geq 0$. Then, following standard notations, we define the $\sin\Theta$-matrix as
$\sin\Theta(V,\hat{V})=\operatorname{diag}\left[\sin(\cos^{-1}\left(\sigma_{1}\right)),\cdots,\sin(\cos^{-1}\left(\sigma_{d}\right))\right],$
and use $\|\sin\Theta(V,\hat{V})\|_{2}$ to measure the similarity between $V$ and $\hat{V}$. Further, for symmetric matrices $Z,\hat{Z}\in\mathbb{R}^{n\times n}$ with leading $d<n$ eigenvectors denoted as $V,\hat{V}\in\mathbb{R}^{n\times d}$, respectively, we define $\sin_{d}(Z,\hat{Z}):=\|\sin\Theta(V,\hat{V})\|_{2}$. This is uniquely defined only if either $|\gamma_{d}|>|\gamma_{d+1}|$ or $|\gamma_{d}|=|\gamma_{d+1}|=0$. For all matrices we use this for, this will be true for sufficiently large $n$.
Lemma 1 in Cai and Zhang (2016) establishes the equivalence between the sin distance and several other metrics. We will use the following two corollaries from this Lemma:
@@ B.3 Section 4 Proofs (Consistency)
Another important result we use is Theorem 1 of Cai and Zhang (2016), which provides a useful bound on the $\sin$-distance between the eigenspaces of a matrix and its perturbation. Since originally this result was formulated for arbitrary matrices and we only need it for symmetric matrices, we restate it in Lemma B.1 with a slightly looser but much more convenient bound.
Proof of Theorem 4.1.
Assume $\mathcal{E}_{noise}$ in (20) and initialization error event in (22) occur simultaneously, which has probability at least $1-(M+K+1)ne^{-C_{0}n}$ by Lemma B.3 and Assumption 3.
The initial step is to rewrite the first iteration update for each parameter in Stage I and Stage II as a soft-thresholding operator applied to the parameter’s ground-truth value, say $Z$, plus the error term, which comprises Gaussian noise and the weighted sum of an update-dependent set of latent components errors:
@@
The first stage update rule for the parameters $S+Q_{k}$ and $\{R_{k\ell}\}_{\ell=1}^{m_{k}}$ can be rewritten
@@
where by $\Delta_{Z}^{(1)}:=Z^{(1)}-Z$ we denote the error of parameter $Z$ after the first iteration. The second stage update rule at the first iteration can be written similarly:
$Q_{k}^{(1)}=\mathcal{T}_{\rho_{2k}}\Bigl[Q_{k}-\Delta_{S}^{(0)}-{1\over m_{k}}\sum_{\ell=1}^{m_{k}}\Delta_{R_{k\ell}}+\bar{E}_{k}\Bigr],\qquad k=1,\ldots,K,$ (36)
$S^{(1)}=\mathcal{T}_{\rho_{2}}\Bigl[S-\sum_{k=1}^{K}{m_{k}\over M}\Delta_{Q_{k}}^{(1)}-{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Delta_{R_{k\ell}}+\bar{E}\Bigr],$
where $\Delta_{R_{k\ell}},(k,\ell)\in\mathcal{I}$ denotes the individual component error after the $k$-th group updates in the first stage.
The key to our proof will be Lemma B.2 that shows that the soft threshold operator applied to a matrix perturbed by additive error as in (34) should have the threshold $\rho$ with the rate of the total error’s spectral norm to guarantee that the estimate $\hat{Z}$ is close to the ground truth $Z$ in the Frobenius norm. In particular, it should dominate both the spectral norm of the other components’ errors, which we control using Lemma B.3, and the average of the appropriate Gaussian noise matrices, which we control by restricting the analysis to the set $\mathcal{E}_{noise}$. In what follows, we formalize this intuition for Stage I and then for Stage II.
Stage I (group $k$): We first establish the properties of $R_{k\ell}^{(1)},\ \ell=1,\ldots,m_{k}$ by applying Lemma B.2 with
@@
and by Property 4, if $R_{k\ell}$ is PSD, then $R_{k\ell}^{(1)}$ is also PSD. Combining Corollary B.1 and (37), we have, for sufficiently large $n$,
@@
which implies together with (37) that,
@@
Next, we establish an error bound for $\Delta_{S+Q_{k}}^{(1)}$. With $C_{2}:={11(1+3\sigma)B_{R}\over b_{R}(1\wedge b_{R})}$, define
$\rho_{1k}=C_{2}\max_{1\leq\ell\leq m_{k}}\rho_{1k\ell}\Bigl[{1\over m_{k}}\vee s_{u,u}^{(k)}\vee n^{-\tau}\max_{1\leq\ell\leq m_{k}}\rho_{1k\ell}\Bigr]^{1/2},$ (41)
so that, since $\rho_{1k\ell}\geq 3\sigma(n/m_{k})^{1/2}$,
@@
Then Properties 1 and 2 in Lemma B.2 imply
@@
Stage II: We first establish the properties of $Q_{k}^{(1)},\ k=1,\ldots,K$ by applying Lemma B.2 with the following correspondence of notations due to (36)
@@
Then, Lemma B.3 with $s:=s_{w,w}$, $\theta:=\max_{1\leq k\leq K}\theta_{k}$, and $\rho:=\max_{1\leq k\leq K}\rho_{2k}$ implies
$\Bigl\|{1\over M}\sum_{k=1}^{K}m_{k}\Delta_{Q_{k}}^{(1)}\Bigr\|_{2}\leq c_{K}\Bigl\|{1\over K}\sum_{k=1}^{K}\Delta_{Q_{k}}^{(1)}\Bigr\|_{2}\leq{11c_{K}B_{Q}\over b_{Q}(1\wedge b_{Q})}\max_{1\leq k\leq K}\rho_{2k}\Bigl[{1\over K}\vee s_{w,w}\vee n^{-\tau}\max_{1\leq k\leq K}\rho_{2k}\Bigr]^{1/2},$
where $c_{K}>0$ is a constant such that $m_{k}/M\leq c_{K}/K$ for each $k=1,\ldots,K$ (it is guaranteed to exist for sufficiently large $n$ by Assumption 1). To bound $\Delta_{S}^{(1)}$, we first apply Lemma B.3 with $s:=s_{u,u},\theta:=\max_{(k,\ell)\in\mathcal{I}}\theta_{k\ell}$, and $\rho:=\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}$ to obtain
$\Bigl\|{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Delta_{R_{k\ell}}\Bigr\|_{2}$ $\leq 11{B_{R}\over b_{R}}\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}\Bigr[{1\over M}\vee s_{u,u}\vee{1\over b_{R}n^{\tau}}\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}\Bigl]^{1/2}.$ (46)
Letting $C_{3}:=12\bigl({c_{K}B_{Q}\over b_{Q}(1\wedge b_{Q})}+{B_{R}\over b_{R}(1\wedge b_{R})}\bigr)$, define
@@
so that by ${\rho_{1k\ell}/M^{1/2}}\geq 3\sigma(n/M)^{1/2}\geq\|\bar{E}\|_{2}$ we have
$\Bigl\|{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Delta_{R_{k\ell}}^{(1)}\Bigr\|_{2}+\Bigl\|{1\over M}\sum_{k=1}^{K}m_{k}\Delta_{Q_{k}}^{(1)}\Bigr\|_{2}+\|\bar{E}\|_{2}<\rho_{2}.$
@@
By Property 4, if $S$ is PSD, then $S^{(1)}$ is also PSD. ∎
We conclude this section by proving Proposition 4.1, establishing the bounds on the first and second stage averaging initializers.
Proof of Proposition 4.1.
We rewrite the first and second stage initializers similarly to (34) by substituting soft thresholding with hard thresholding:
$(S+Q_{k})^{(0)}=\Bigl[{1\over m_{k}}\sum_{\ell=1}^{m_{k}}A_{k\ell}\Bigr]_{d_{0}+d_{k}}=\Bigl[(S+Q_{k})-{1\over m_{k}}\sum_{\ell=1}^{m_{k}}R_{k\ell}+\bar{E}_{k}\Bigr]_{d_{0}+d_{k}},$
$S^{(0)}=\Bigl[{1\over K}\sum_{k=1}^{K}{1\over m_{k}}\sum_{\ell=1}^{m_{k}}(A_{k\ell}-\hat{R}_{k\ell})\Bigr]_{d_{0}}=\Bigl[S+{1\over K}\sum_{k=1}^{K}Q_{k}-{1\over K}\sum_{k=1}^{K}{1\over m_{k}}\sum_{\ell=1}^{m_{k}}\Delta_{R_{k\ell}}+\bar{E}_{k}\Bigr]_{d_{0}+d_{k}}.$
Our next goal is to use Lemma B.4 to establish deterministic bounds on the errors of two initializers, assuming $\mathcal{E}_{noise}$ holds.
Stage I. By Lemma 4 in (MacDonald et al., 2022), we have
$\|{1\over m_{k}}\sum_{\ell=1}^{m_{k}}R_{k\ell}\|_{2}\leq B_{R}n^{\tau}(m_{k}^{-1}+s_{u,u}^{(k)})^{1/2}$
Therefore, by triangular inequality, it holds for $E:=\bar{E}_{k}-{1\over m_{k}}\sum_{\ell=1}^{m_{k}}R_{k\ell}$:
$\|E\|_{2}\leq 3\sigma(n/m_{k})^{1/2}+B_{R}n^{\tau}\bigl(m_{k}^{-1}+s_{u,u}^{(k)}\bigr)^{1/2}=o(n^{\tau})$
and by Property 2 in Lemma B.4, we obtain for sufficiently large $n$
$\|(S+Q_{k})-(S+Q_{k})^{(0)}\|_{2}\leq{19B_{S+Q}\over b_{S+Q}}\|\mathcal{P}_{[V,W_{k}]}E\|_{2}.$
The needed result follows by triangular inequality and Lemma 4 in (MacDonald et al., 2022):
$\|\mathcal{P}_{[V,W_{k}]}E\|_{2}$ $\leq\|\bar{E}_{k}\|_{2}+\|\sum_{\ell=1}^{m_{k}}\mathcal{P}_{[V,W_{k}]}R_{k\ell}\|_{2}/m_{k}$
$\leq 3\sigma(n/m_{k})^{1/2}+B_{R}n^{\tau}s_{vw,u}^{(k)}(m_{k}^{-1}+s_{u,u}^{(k)})^{1/2}\lesssim n^{1/2}.$
Stage II. Combining (46) with the established rates $r^{(I)}_{k}\lesssim n^{1/2}$ for the first-stage initializers above, we obtain for sufficiently large $n$,
$\Bigl\|{1\over K}\sum_{k=1}^{K}{1\over m_{k}}\sum_{\ell=1}^{m_{k}}\Delta_{R_{k\ell}}\Bigl\|_{2}\leq c_{K}^{\prime}\Bigl\|{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Delta_{R_{k\ell}}\Bigr\|_{2}\lesssim(n/M)^{1/2},$
where $c_{K}^{\prime}>0$ is a constant such that $M/K\leq c_{K}^{\prime}m_{k}$ for every $k=1,\ldots,K$ (it is guaranteed to exist by Assumption 1). By Lemma 4 in (MacDonald et al., 2022), we also have
$\Bigl\|{1\over K}\sum_{k=1}^{K}Q_{k}\Bigr\|_{2}\leq B_{Q}n^{\tau}(K^{-1}+s_{w,w})^{1/2}$
Therefore, by triangular inequality, it holds for $E:={1\over K}\sum_{k=1}^{K}Q_{k}+\bar{E}-{1\over K}\sum_{k=1}^{K}{1\over m_{k}}\sum_{\ell=1}^{m_{k}}\Delta_{R_{k\ell}}$ if $n$ is sufficiently large
$\|E\|_{2}\leq(1+\delta)B_{Q}n^{\tau}(K^{-1}+s_{w,w})^{1/2}.$
This implies $|\gamma_{d_{0}}(S)|\geq b_{S}n^{\tau}\geq 4\|E\|_{2}$ by (25). So, we can apply Property 2 in Lemma B.4 to obtain:
$\|S^{(0)}-S\|_{2}\leq{19B_{S}\over b_{S}}\|\mathcal{P}_{V}E\|_{2}.$
Finally, Lemma 4 in (MacDonald et al., 2022) and triangular inequality imply the needed rate
$\|\mathcal{P}_{V}E\|_{2}$ $\leq\|\sum_{k=1}^{K}\mathcal{P}_{V}Q_{k}\|_{2}/K+O(\sqrt{n/M})$
$\leq B_{Q}n^{\tau}s_{v,w}(K^{-1}+s_{w,w})^{1/2}+O(\sqrt{n/M})\lesssim\sqrt{nK/M}.$